# FaceShield: Defending Facial Image against Deepfake Threats

### v2

@@ E Evaluating FaceShield under Image Purifications
These experiments are conducted using the IP-Adapter model [61], with the same dataset as in Table 1. The quantitative results for ISM and PSNR are presented in Fig.14, while the qualitative results are shown in Fig.15 and Fig.16. As shown in the results, FaceShield causes only minor performance degradation across various purification methods, yet still demonstrates superior performance compared to other baselines [32, 31, 42, 60], proving its remarkable robustness.
Figure 14: Quantitative results of FaceShield-protected images after passing through various purification methods and evaluated on a deepfake model [61]. Our method demonstrates robustness against various purification methods, including JPEG compression, bit reduction, and two types of resizing, with its performance compared to baseline methods [32, 31, 42, 60]. The results, measured using PSNR and Identity Score Matching (ISM), show that our method closely resembles lossless (PNG) outcomes while consistently outperforming the baselines. Both metrics indicate better performance with lower values.
@@ F Additional Qualitative Results
In this section, we present additional qualitative results of our methods. Specifically, Fig.17 to Fig.19 compare our approach with baseline methods [32, 31, 42, 60] on various diffusion-based deepfake models [25, 68, 50, 61], using a pair of source and target images. Fig.20 compares our method with the baselines on the FaceSwap via Diffusion model [50] across different image pairs. Fig.21 shows the comparison within the IP-Adapter model [61], while Fig.22 compares our method with the baselines on the DiffSwap model [68]. Fig.23 presents a comparison on the DiffFace model [25]. Finally, Fig.24 and Fig.25 showcase additional experiments on two GAN-based deepfake models: SimSwap [4] and InfoSwap [14], respectively.
Figure 15: The results of applying three levels of JPEG compression and two levels of bit reduction to images protected by FaceShield, followed by evaluation on a deepfake model [61], show that the performance degradation is minimal compared to lossless storage (PNG).
Figure 16: The results of applying two types of resizing methods, with 75% and 50% scaling, to images protected by FaceShield, followed by evaluation on a deepfake model [61], show that the performance degradation is minimal compared to lossless storage (PNG).
Figure 17: Qualitative comparisons with AdvDM [32], Mist [31], PhotoGuard [42], and SDST [60] across four diffusion-based deepfake models: FaceSwap [50], IP-Adapter [61], DiffSwap [68], and DiffFace [25].
Figure 18: Qualitative comparisons with AdvDM [32], Mist [31], PhotoGuard [42], and SDST [60] across four diffusion-based deepfake models: FaceSwap [50], IP-Adapter [61], DiffSwap [68], and DiffFace [25].
Figure 19: Qualitative comparisons with AdvDM [32], Mist [31], PhotoGuard [42], and SDST [60] across four diffusion-based deepfake models: FaceSwap [50], IP-Adapter [61], DiffSwap [68], and DiffFace [25].
Figure 20: Qualitative comparisons for FaceSwap [50].
Figure 21: Qualitative comparisons for IP-Adapter [61].
Figure 22: Qualitative comparisons for DiffSwap [68].
Figure 23: Qualitative comparisons for DiffFace [25].
Figure 24: Qualitative results for SimSwap [4].
Figure 25: Qualitative results for InfoSwap [14].
@@ G Additional Experiments
IP-Adapter [61] is a lightweight adapter that enables image conditions in pre-trained text-to-image diffusion models [39]. Previous approaches [41, 26] that utilized image conditions primarily relied on fine-tuning text-conditioned diffusion models. However, these methods often demanded significant computational resources and resulted in models that were challenging to reuse. To address these limitations, the IP-Adapter, which proposes a decoupled cross-attention mechanism, has drawn considerable attention for its practical applicability. It is commonly used in inpainting methods with image conditions. As shown in Fig.26, multiple versions of the IP-Adapter model have been developed with Stable diffusion v1.5 [39].
A more detailed look at the various models reveals that the original model [61] uses the CLIP image encoder [38] to extract features from the input image. In contrast, the IP-AdapterXL improves on this by utilizing larger image encoders, such as ViT-BigG or ViT-H, which enhance both capacity and performance. On the other hand, the IP-AdapterPlus and XL versions modify the architecture by adopting a patch embedding method inspired by Flamingo’s perceiver resampler [2], allowing for more efficient image encoding. Similarly, the IP-AdapterFaceID and XL versions replace the CLIP image encoder with InsightFace, extracting FaceID embeddings from reference images. This enables the combination of additional text-based conditions with the facial features of the input image, allowing for the generation of diverse styles. The IP-AdapterFaceIDPlus and XL versions further enhance the image encoding pipeline by incorporating multiple components. InsightFace is used for detailed facial features, the CLIP image encoder captures global facial characteristics, and the Perceiver-resampler effectively combines these features to improve the model’s overall functionality.
Qualitative results. We evaluate the transferability across different IP-Adapter versions and present comparisons with baseline methods. Specifically, we conducted experiments on eight of these models, with results and model descriptions provided in Fig.27 to Fig.30. These results demonstrate the versatility of FaceShield, showing that it is applicable across various sub-models of the IP-Adapter [61].

### v3

@@ E Evaluating FaceShield under Image Purifications
These experiments are conducted using the IP-Adapter model [ye2023ip], with the same dataset as in Table 1 in the main paper. The quantitative results for ISM and PSNR are presented in Fig.14, while the qualitative results are shown in Fig.15 and Fig.16. As shown in the results, FaceShield causes only minor performance degradation across various purification methods, yet still demonstrates superior performance compared to other baselines [liang2023adversarial, liang2023mist, salman2023raising, xue2023toward], proving its remarkable robustness.
Figure 14: Quantitative results of FaceShield-protected images after passing through various purification methods and evaluated on a deepfake model [ye2023ip]. Our method demonstrates robustness against various purification methods, including JPEG compression, bit reduction, and two types of resizing, with its performance compared to baseline methods [liang2023adversarial, liang2023mist, salman2023raising, xue2023toward]. The results, measured using PSNR and Identity Score Matching (ISM), show that our method closely resembles lossless (PNG) outcomes while consistently outperforming the baselines. Both metrics indicate better performance with lower values.
@@ F Additional Qualitative Results
In this section, we present additional qualitative results of our methods. Specifically, Fig.17 to Fig.19 compare our approach with baseline methods [liang2023adversarial, liang2023mist, salman2023raising, xue2023toward] on various diffusion-based deepfake models [kim2212diffface, zhao2023diffswap, wang2024face, ye2023ip], using a pair of source and target images. Fig.20 compares our method with the baselines on the FaceSwap via Diffusion model [wang2024face] across different image pairs. Fig.21 shows the comparison within the IP-Adapter model [ye2023ip], while Fig.22 compares our method with the baselines on the DiffSwap model [zhao2023diffswap]. Fig.23 presents a comparison on the DiffFace model [kim2212diffface]. Finally, Fig.24 and Fig.25 showcase additional experiments on two GAN-based deepfake models: SimSwap [chen2020simswap] and InfoSwap [gao2021information], respectively.
Figure 15: The results of applying three levels of JPEG compression and two levels of bit reduction to images protected by FaceShield, followed by evaluation on a deepfake model [ye2023ip], show that the performance degradation is minimal compared to lossless storage (PNG).
Figure 16: The results of applying two types of resizing methods, with 75% and 50% scaling, to images protected by FaceShield, followed by evaluation on a deepfake model [ye2023ip], show that the performance degradation is minimal compared to lossless storage (PNG).
Figure 17: Qualitative comparisons with AdvDM [liang2023adversarial], Mist [liang2023mist], PhotoGuard [salman2023raising], and SDST [xue2023toward] across four diffusion-based deepfake models: FaceSwap [wang2024face], IP-Adapter [ye2023ip], DiffSwap [zhao2023diffswap], and DiffFace [kim2212diffface].
Figure 18: Qualitative comparisons with AdvDM [liang2023adversarial], Mist [liang2023mist], PhotoGuard [salman2023raising], and SDST [xue2023toward] across four diffusion-based deepfake models: FaceSwap [wang2024face], IP-Adapter [ye2023ip], DiffSwap [zhao2023diffswap], and DiffFace [kim2212diffface].
Figure 19: Qualitative comparisons with AdvDM [liang2023adversarial], Mist [liang2023mist], PhotoGuard [salman2023raising], and SDST [xue2023toward] across four diffusion-based deepfake models: FaceSwap [wang2024face], IP-Adapter [ye2023ip], DiffSwap [zhao2023diffswap], and DiffFace [kim2212diffface].
Figure 20: Qualitative comparisons for FaceSwap [wang2024face].
Figure 21: Qualitative comparisons for IP-Adapter [ye2023ip].
Figure 22: Qualitative comparisons for DiffSwap [zhao2023diffswap].
Figure 23: Qualitative comparisons for DiffFace [kim2212diffface].
Figure 24: Qualitative results for SimSwap [chen2020simswap].
Figure 25: Qualitative results for InfoSwap [gao2021information].
@@ G Additional Experiments
IP-Adapter [ye2023ip] is a lightweight adapter that enables image conditions in pre-trained text-to-image diffusion models [Rombach_2022_CVPR]. Previous approaches [ruiz2023dreambooth, kumari2023multi] that utilized image conditions primarily relied on fine-tuning text-conditioned diffusion models. However, these methods often demanded significant computational resources and resulted in models that were challenging to reuse. To address these limitations, the IP-Adapter, which proposes a decoupled cross-attention mechanism, has drawn considerable attention for its practical applicability. It is commonly used in inpainting methods with image conditions. As shown in Fig.26, multiple versions of the IP-Adapter model have been developed with Stable diffusion v1.5 [Rombach_2022_CVPR].
A more detailed look at the various models reveals that the original model [ye2023ip] uses the CLIP image encoder [radford2021learning] to extract features from the input image. In contrast, the IP-AdapterXL improves on this by utilizing larger image encoders, such as ViT-BigG or ViT-H, which enhance both capacity and performance. On the other hand, the IP-AdapterPlus and XL versions modify the architecture by adopting a patch embedding method inspired by Flamingo’s perceiver resampler [alayrac2022flamingo], allowing for more efficient image encoding. Similarly, the IP-AdapterFaceID and XL versions replace the CLIP image encoder with InsightFace, extracting FaceID embeddings from reference images. This enables the combination of additional text-based conditions with the facial features of the input image, allowing for the generation of diverse styles. The IP-AdapterFaceIDPlus and XL versions further enhance the image encoding pipeline by incorporating multiple components. InsightFace is used for detailed facial features, the CLIP image encoder captures global facial characteristics, and the Perceiver-resampler effectively combines these features to improve the model’s overall functionality.
Qualitative results. We evaluate the transferability across different IP-Adapter versions and present comparisons with baseline methods. Specifically, we conducted experiments on eight of these models, with results and model descriptions provided in Fig.27 to Fig.30. These results demonstrate the versatility of FaceShield, showing that it is applicable across various sub-models of the IP-Adapter [ye2023ip].